1 pound of peanuts cost $2.80 and 1 pound of cashews cost $5.30. 1 pound of mixed nuts cost $3.30. How many pounds of peanuts does it take to make 100 pounds of mixed nuts?
Answer by mananth(16946)

---------price ---------------- quantity
Cashews 5.3 ---------------- x pounds
peanuts 2.8 ------ 100 - x pounds
Mixture 3.30 ---------------- 100

5.3 x + 2.8 ( 100 - x ) = 3.3 * 100

5.3x+ 280-2.8x =330
5.3x-2.8x=330-280
2.5x=50
/ 2.5
x=20 pounds $5.30 Cashews
80 pounds $2.80 peanuts